# Explainable Deep Learning Analysis for Raga Identification in Indian Art Music

### v1

@@ Abstract
The task of Raga Identification is a very popular research problem in Music Information Retrieval. Few studies that have explored this task employed various approaches, such as signal processing, Machine Learning (ML) methods, and more recently Deep Learning (DL) based methods. However, a key question remains unanswered in all of these works: do these ML/DL methods learn and interpret Ragas in a manner similar to human experts? Besides, a significant roadblock in this research is the unavailability of ample supply of rich, labeled datasets, which drives these ML/DL based methods. In this paper, we introduce ”Prasarbharti Indian Music” version-1 (PIM-v1), a novel dataset comprising of 191 hours of meticulously labeled Hindustani Classical Music (HCM) recordings, which is the largest labeled dataset for HCM recordings to the best of our knowledge. Our approach involves conducting ablation studies to find the benchmark classification model for Automatic Raga Identification (ARI) using PIM-v1 dataset. We achieve a chunk-wise f1-score of 0.89 for a subset of 12 Raga classes. Subsequently, we employ model explainability techniques to evaluate the classifier’s predictions, aiming to ascertain whether they align with human understanding of Ragas or are driven by arbitrary patterns. We validate the correctness of model’s predictions by comparing the explanations given by two ExAI models with human expert annotations. Following this, we analyze explanations for individual test examples to understand the role of regions highlighted by explanations in correct or incorrect predictions made by the model.
@@ I Introduction
Indian Art music (IAM) is a culturally rich and complex musical tradition that has evolved over centuries. It is categorized into two main categories: Hindustani Classical Music (HCM) and Carnatic music (CM). HCM is predominant in the northern regions of India, while Carnatic music is more prevalent in the south. Both forms of music have their own distinct nature and way of singing. Central to IAM is the concept of a Raga, which can be termed as melodic structures that provide a framework for improvisation and composition, making it fundamental descriptor for IAM. It represents the mood of any musical composition. Ragas also form the basis of a lot of indian film or folk music[1]. Detailed description about IAM Ragas can be found at [2, 3].
To compare with western music, in HCM, a raga is not merely defined by its set of notes, but also by the way these notes are performed, with emphasis on smooth melodic glides and the space between notes. This contrasts with Eurocentric music genres, where a scale is primarily defined by its set of notes[4]. Also, in western music, there is more emphasis on harmony, whereas HCM focuses on melody[5]. Raga is also different from a mode in Eurocentric music as it encompasses a broader range of musical features including scale, notes’ sequence, hierarchy, relative duration, micro-tonal intonation, ornamentation, melodic movement, specific phrases, and mood[4]. Due to these distinctions, the task of Automatic ARI in IAM differs significantly from genre classification in Western music. In IAM, categories such as Dhrupad, Khayal, Thumri, Dadra, Tappa, and Bhajan etc can be considered distinct genres[5], but not Ragas.
IAM’s emphasis on improvisation within the framework of Raga underscores the intricate nature of this musical tradition, setting it apart from Western music genres.
For data-driven methods used in computational musicology and music information retrieval (MIR), digital music artifacts with metadata and expert annotations are a prerequisite[6]. Today, searching for music based on identifiers like ragas is not as simple as searching for text on search engines like Google. Efforts are underway to automate this process, starting with the automatic identification of ragas. Music enthusiasts and researchers who view music as more than just entertainment require a music/audio retrieval system based on melody or raga[7]. ARI has several other applications, including music recommendation, music teaching, and automatic annotation.
@@
The progress of research in IAM is often hindered by the scarcity of a variety of open labeled datasets. Introduction of new datasets in IAM can be useful for making tasks like melody extraction, ARI, tala identification, source separation, automatic mixing, and performance analysis more robust and generalisable.
@@
For the task of ARI, various Machine Learning(ML)/deep learning(DL) based approaches are being used, but we have no way to know if the models are making predictions as if an IAM expert would do or not. This motivates us to look for these answers from the field of Explainable AI (ExAI), which has barely been explored in the field of IAM analysis. ExAI is a popular research area within ML that aims to provide explanations about complex deep learning models, making them interpretable. ExAI focuses on explaining the behavior of models and their predictions, in various known and unknown situations, in a post-hoc manner, hence increasing transparency and understanding of these models, ultimately improving their reliability and trustworthiness. In this study, we take up the task of ARI for IAM and then use ExAI methods to show the reliability of the classifier.
@@
  • We Introduce a novel HCM dataset along with manual annotations, and benchmark the dataset for the task of Raga classification by undergoing Ablation Studies. We release this dataset in the form of 30-second audio chunks along with labels and additional metadata. The sample audios can be accesed at: https://github.com/ParampreetSingh97/PIM_v1_ExAI.git.
  • We make one of the first attempts on using ExAI techniques for raga identification, and subsequently validating the significance and correctness of the Raga classifier musically.
@@ II-A HCM Datasets
The Indian Art Music Raga Recognition Dataset (RRD)[8], offers 116 hours of HCM with 300 recordings featuring 30 unique ragas, with 10 recordings per raga. Then there is Saraga dataset [9], which comprises of 108 recordings of HCM in 61 unique Ragas and 9 unique talas, totaling 43.6 hours. Both of these datasets feature live performances by various experts, along with providing labels like Raga names, manually marked melodic phrases, and features like tonic and pitch extracted using various algorithms. While Saraga allows download access with annotations, RRD requires access requests for audio file viewing. Notably, we observe a significant overlap of audio files between Saraga and RRD. Some other works [10, 11] utilize self-made HCM datasets for ARI, but these datasets have not been made publicly available.
@@ II-B Raga Identification
For HCM, the paper [10] uses their own generated instrumental dataset of 20 second snippets for 12 Raga classification. They make use of a combination of features like Pitch, MFCC and other statistics derived from mfcc together for the classification task. Another paper [4] does 2-Raga classification using a subset from Saraga Dataset and report that using Swargram features is helpful for Raga classification, which is nothing but Chromagram features shifted in pitch. The paper [12] uses pitch class distribution and N-gram approach on the extracted pitch for raga identification and reports that longer audio clip length is beneficial for the task. Another work [13] uses a dataset downloaded from youtube for the task of 6-class classification. In paper [14], KNN and SVM machine learning algorithms were used for classifying different ragas like Yaman and Bhairavi, and the KNN was found to be performing slightly better. The paper [15] proposed a hierarchical deep learning system called PhonoNet, which used a CNN model in the first stage to predict the raga of a small audio chunk, and a two-stage recurrent LSTM network in the second stage to look at various parts of the audio input and prioritize observed patterns.
@@
There are some other works related to Carnatic Music[1, 16, 17, 18, 19, 20, 21, 22], and most of them use basic features like extracted pitch, MFCCs, Spectral centroid, chromagram, spectrogram, scalogram etc and apply different ML classifiers like Naive Bayes, SVM, Random forest, GMMs, and some also use DL models like RNN, LSTM, CNN etc.
@@
We observe that we have more research centered around Carnatic music, primarily because of higher availability of labeled datasets as compared to HCM. Also, to the best of our knowledge, none of the works in the literature have explored ExAI methods for the task of Raga Identification in any form.
II-C ExAI methods
@@
Until now, we have a lot of research already available focused on ExAI methods for DL models used for various tasks involving Images, and some works for audio and music also. We will now discuss some notable works in both. One very prominent kind of approach for post-hoc explainability involves the use of model’s Gradients from various layers of a CNN or ANN model and using them in a intelligent way to visualize explanations of the classifier in the form of saliency map, superimposed over the original image itself, highlighting the most important region for classification by the classifier [23, 24, 25, 26, 27, 28, 29, 30]. We will be using GradCAM++[28], which is a very popular method for CNN explanations. It weighs each of the pixels in each feature map of the last CNN layer separately, with weights determined using the gradients of these layers in a clever way, and then finally stacking all the weighted feature maps together to get the final attribution map, which highlights important regions in an image. In the audio domain, we have some works  [31, 32] which treat audio feature representations as images and use the ExAI methods for images to generate explanations.
@@
Other type is of black box models, whose gradients are not accessible and we have to design ExAI methods by querying the model multiple times. There are some notable works for black box model explainability, with LIME[33] generating locally faithful explanations by training an interpretable model on perturbed instances of the input data sample. Another method SHAP[34], uses game theory to assign a unique contribution score to each feature in the input, indicating its impact on the model’s output. SHAP is computationally intensive, especially for models with a large number of features. Another approach Anchors[35] express explanations as simple IF-THEN rules, making them more easily understandable, but it struggles with large and complex datasets.
@@
For audio , we have a variant of LIME in SoundLIME[36], which works on spectrograms or other audio based features, treats them as image data only and then replicate the LIME framework.
For our work, we employ a very popular white-box ExAI model in Grad-CAM++[28], and we use SoundLIME out of the black box models. We employ both white-box and black-box ExAI models to ensure comprehensive coverage of interpretability techniques, enabling a thorough comparison and analysis of the performance of our ARI classifier model.
@@ III Dataset
The dataset comprises a total of 501 HCM songs with each song being a polyphonic audio containing vocals of the main artist, percussion instruments like tabla, Pakhawaj, along with Manjeera in some audios, along with non-percussion instruments like Harmonium, Sarangi, Violin, and also tanpura (drone) which provides the tonic for the whole music performance. These music recordings are sung in variety of Ragas and Talas with all of the songs being of varying lengths. The dataset includes 144 unique Ragas performed by 169 vocal artists. The Distribution of Ragas is shown in table: I. Harmonium is the most common instrument, present in 372 songs, while some performances feature multiple string instruments. All of the songs include drone (tanpura) playing in the background of the performance. The dataset includes performances in 20 different talas accompnying the vocals, with teentaal (121 instances) and ektaal (146 instances) being the most frequently used talas as shown in fig:1. Each song is available in the form of 30-second chunks, totaling 23,005 audio files, approximately 191.70 hours of music.
@@ III-B Recording Source and Quality
All recordings are sourced from the archives of Prasar Bharti (Akashwani, Delhi), which are not available online elsewhere. They feature performances by HCM ustaads (experts) and are mainly studio recordings or recordings from live shows, ensuring high audio quality. Most of the featured vocal, percussion, or string instrument artists are well-known in the HCM community. All audio files are mono tracks. Some of the recordings feature speech at the start which includes introduction of the artists and some details about their performance and ranges from a few seconds to about a few minutes in some recordings.
@@ III-C Metadata
The metadata contains details such as the names of vocal and instrumental artists, the instruments played, the event where the performance took place, the name of the song, and the number of 30-second files for each song. We also add to the metadata the Raga name, the tonic and the start and end times of music and speech for each of those songs, labeled manually by our annotators.
@@ III-D Manual Annotation Process
Two music experts with over 10 years of experience in vocal music teaching and singing for HCM carried out the task of annotation. They identified the Raga and tonic of each of the 501 music performances, listed in the metadata. This annotation task involves labelling the whole audio with a single Raga and Tonic label, as they do not change throughout a single HCM performance. As the task does not involve frame-wise labelling or subjective labelling, there is negligible chance of discrepancies among annotators, so inter-annotator agreement is not necessary.
TABLE I: Distribution of number of audio files per Raga
No. of files Ragas
1 Abhogi_kanada, Alhaiya_bilawal, Asavari, Asavari (Komal_rishabh), Basant_bhairav, Bhimpalasi, Bhoop_nat, Dagori, Deepak, Des_and_tilak_kamod_combination, Deshi, Desi_todi, Devgandhar, Devgiri_bilawal, Gaud_sarang, Gunkali, Jhoola, Kafi_malhar, Kans_gandhar, Kaushiki_kanada, Lalit, Madhmad_sarang, Malgunji, Mian_ki_todi, Mishra_gara, Mishra_jhinjhoti, Mishra_jinjhoti, Mishra_kirwani, Nat_bhairav, Nat_bihag, Nat_kedar, Pancham, Prabhat_bhairav, Purvi, Ramkali, Sahna_kanada, Salagwarali, Shahana_kanada, Shiv_kalyan, Shivranjani, Shuddha_bairav, Shuddha_basant, Shuddha_sarang, Sohni (pancham), Suha_malhar, Tilak_shyam, Todi, Todi (ahiri), Vrindavani_sarang
2 Anand_bhairav, Bageshri_bahar, Bahar, Bhairav, Bhatiyar, Bihari, Chandrakauns (Bageshri_ang), Chandrakauns (Bhairavi_ang), Charukeshi, Darbari, Deshkaar, Hemant, Hindol, Jait_kalyan, Jayant_malhar, Kirwani, Marwa, Paraj, Paraj_kalingda, Pilu, Puriya_dhanashri, Raysa_kanada, Sampoorna_malkauns, Shahana, Shivmat_bhairav, Swanandi, Tilak_kamod
3 Abhogi, Bageshri_kanada, Basant_bahar, Basant_kedar, Bhinna_shadaja, Durga, Gavati, Gujri_todi, Hameer, Hansdhwani, Janasammohini, Jhinjhoti, Kafi_kanada, Kaunsi_kanada, Mishra_bhairavi, Mishra_desh, Mishra_pahadi, Mishra_tilak_kamod, Mishra_tilang, Sawani_bihag, Shyam_kalyan, Tilang
4 Aheer_bhairav, Basant, Bihagda, Gaud_malhar, Jaunpuri, Kamod, Kaushik_dhwani, Majh_khamaj, Megh, Nand, Rageshree, Saraswati, Surdasi_malhar
5 Adana, Gorakh_kalyan, Madhukauns, Mishra_pilu, Puriya
6 Chayanat, Darbari_kanada, Jaijaiwanti, Kalavati, Mian_malhar, Shankara
8 Chandrakauns, Des, Kafi, Pahadi
9 Kedar, Mishra_khamaj, Puriya_kalyan
10 Jog_kauns, Yaman
11 Malkauns, Maru_bihaag
12 Bageshri, Bhopali, Bihag, Nayaki_kanada, Shuddha_kalyan
13 Khamaj
14 Jog
15 Bhairavi, Sohni
@@ IV Methodology
In this study, we use our PIM dataset to extract chromagram features from the audios and train a CNN-LSTM model to perform multi-class classification for 12 Raga classes. Next, we employ Grad-CAM++ (GC) and SoundLIME (SL), two popular ExAI models, for saliency visualisation and comparing them with annotations from music experts. The sub-sections below explain the entire procedure.

### v2

@@ Abstract
Raga identification is an important problem within the domain of Indian Art music, as Ragas are fundamental to its composition and performance, playing a crucial role in music retrieval, preservation, and education. Few studies that have explored this task employ approaches such as signal processing, Machine Learning (ML), and more recently, Deep Learning (DL) based methods. However, a key question remains unanswered in all these works: do these ML/DL methods learn and interpret Ragas in a manner similar to human experts? Besides, a significant roadblock in this research is the unavailability of an ample supply of rich, labeled datasets, which drives these ML/DL-based methods. In this paper, firstly we curate a dataset comprising 191 hours of Hindustani Classical Music (HCM) recordings, annotate it for Raga and tonic labels, and train a CNN-LSTM model for the task of Automatic Raga Identification (ARI). We achieve a chunk-wise f1-measure of 0.89 for a subset of 12 Raga classes. Following this, we make one of the first attempts to employ model explainability techniques: SoundLIME and GradCAM++ for Raga identification, to evaluate whether the classifier’s predictions align with human understanding of Ragas. We compare the generated explanations with human expert annotations and further analyze individual test examples to understand the role of regions highlighted by explanations in making correct or incorrect predictions made by the model. Our results demonstrate a significant alignment of the model’s understanding with human understanding, and the thorough analysis validates the effectiveness of our approach.
@@ I Introduction
Indian Art Music (IAM) is a culturally rich and complex musical tradition that has evolved over centuries. It is categorized into two main categories: Hindustani Classical Music (HCM) and Carnatic music (CM). Both forms of music have their own distinct nature and way of singing. Central to IAM is the concept of a Raga, which can be termed as melodic structures that provide a framework for improvisation and composition, making it fundamental descriptor for IAM. It represents the mood of any musical composition. Ragas also form the basis of a lot of indian film or folk music[1]. Detailed description about IAM Ragas can be found at [2, 3].
Comparing with Western Music (WM), a Raga in IAM is more complex than a mode in WM, as it encompasses a broader range of musical features, including a specific scale, characteristic note sequences, hierarchical relationships among notes, relative durations, microtonal intonation, ornamentation (gamaka), melodic movements, specific phrases (pakad), and the evocation of particular moods or emotions[4, 2]. In WM, we see more emphasis on harmony as compared to HCM, which focuses on melody[5] with improvisation within the framework of ragas. The task of Raga identification in IAM differs significantly from key or mode detection in WM, as key detection involves identifying the tonal center and scale of a piece[6], whereas Raga identification requires recognizing complex melodic structures, ornamentations, and expressive nuances that are unique to each Raga. Apart from that, categories such as Dhrupad, Khayal, Thumri, Dadra, Tappa, and Bhajan, etc, in IAM are considered distinct genres [5], representing different styles or forms of musical performance, and ragas serve as the foundational melodic frameworks within these genres.
@@
For data-driven methods used in computational musicology and music information retrieval (MIR), digital music artifacts with metadata and expert annotations are a prerequisite[7]. The progress of research in IAM is often hindered by the scarcity of a variety of open, labeled datasets. The introduction of new datasets in IAM can be useful for making tasks like melody extraction, ARI, tala identification, source separation, automatic mixing, and performance analysis more robust and generalisable.
@@
For the task of Automatic Raga Identification (ARI), various Machine Learning (ML) / Deep Learning (DL) based approaches are being used, but we have no way of knowing if the models are making predictions as if an IAM expert would do or not. This motivates us to look for these answers from the field of Explainable AI (XAI), which has barely been explored in the field of IAM analysis. XAI is a popular research area within ML that aims to provide explanations about complex deep learning models, making them interpretable [8, 9, 10]. XAI focuses on explaining the behavior of models and their predictions in various known and unknown situations in a post-hoc manner, hence increasing transparency and understanding of these models, ultimately improving their reliability and trustworthiness. In this study, we take up the task of ARI for IAM and then use XAI methods to show the reliability of the classifier.
@@
  • We manually annotate an HCM dataset using audio recordings from Prasar Bharti. The sample audios, along with metadata information, can be accessed at: https://github.com/ParampreetSingh97/PIM_v1_ExAI.git.
  • We train a CNN-LSTM-based model for the task of ARI and benchmark its performance on the annotated dataset.
  • We make one of the first attempts at using XAI techniques for ARI and subsequently validating the significance and correctness of the Raga classifier musically.
@@ II-A HCM datasets
The Indian Art Music Raga Recognition Dataset (RRD)[11], offers 116 hours of HCM with 300 recordings featuring 30 unique Ragas, with 10 recordings per Raga. Then there is Saraga dataset [12], which comprises of 108 recordings of HCM in 61 unique Ragas and 9 unique talas, totaling 43.6 hours. Both of these datasets feature live performances by various experts, along with providing labels like Raga names, manually marked melodic phrases, and features like tonic and pitch extracted using various algorithms. While Saraga allows download access with annotations, audio files RRD are currently unavailable. Notably, we observe a significant overlap of audio files between Saraga and RRD. Some other works [13, 14] utilize self-made HCM datasets for ARI, but these datasets have not been made publicly available.
@@ II-B Raga Identification
For HCM, the paper [13] uses their own generated instrumental dataset of 20 second snippets for 12 Raga classification. They make use of a combination of features like Pitch, MFCC and other statistics derived from mfcc together for the classification task. Another paper [4] does 2-Raga classification using a subset from Saraga dataset and report that using Swargram features is helpful for Raga classification, which is nothing but chromagram features shifted in pitch. The paper [15] uses pitch class distribution and N-gram approach on the extracted pitch for ARI and reports that longer audio clip length is beneficial for the task. Another work [16] uses a dataset downloaded from YouTube for the task of 6-class classification. In paper [17], KNN and SVM machine learning algorithms were used for classifying different Ragas like Yaman and Bhairavi, and the KNN was found to be performing slightly better. The paper [18] proposed a hierarchical deep learning system called PhonoNet, which used a CNN model in the first stage to predict the Raga of a small audio chunk, and a two-stage recurrent LSTM network in the second stage to look at various parts of the audio input and prioritize observed patterns.
@@
There are some other works related to CM[1, 19, 20, 21, 22, 23, 24, 25], and most of them use basic features like extracted pitch, MFCCs, spectral centroid, chromagram, spectrogram, scalogram etc and apply different ML classifiers like Naive Bayes, SVM, Random forest, GMMs, and some also use DL models like RNN, LSTM, CNN etc.
@@
We observe that we have more research centered around CM, primarily because of higher availability of labeled datasets as compared to HCM. Also, to the best of our knowledge, none of the works in the literature have explored XAI methods for the task of ARI in any form.
II-C XAI methods
@@
Until now, we have a lot of research already available focused on XAI methods for DL models used for various tasks involving images, and some works for audio and music also. One very prominent kind of approach for post-hoc explainability involves the use of model’s gradients from various layers of a CNN or ANN model and using them in an intelligent way to visualize explanations of the classifier in the form of saliency map superimposed over the original image itself, highlighting the most important region for classification by the classifier [26, 27, 28, 29, 8, 9, 30, 31]. We will be using GradCAM++[9], which is a very popular method for CNN explanations. It weighs each of the pixels in each feature map of the last CNN layer separately, with weights determined using the gradients of these layers, and then finally stacking all the weighted feature maps together to get the final attribution map, which highlights important regions in an image. In the audio domain, we have some works[32] which treat audio feature representations as images and use the XAI methods for images to generate explanations.
@@
Other type is of black box models, whose gradients are not accessible and we have to design XAI methods by querying the model multiple times. There are some notable works for black box model explainability, with LIME[33] generating locally faithful explanations by training an interpretable model on perturbed instances of the input data sample. Another method SHAP[34], uses game theory to assign a unique contribution score to each feature in the input, indicating its impact on the model’s output. SHAP is computationally intensive, especially for models with a large number of features. Another approach Anchors[35] express explanations as simple IF-THEN rules, making them more easily understandable, but it struggles with large and complex datasets.
@@
For audio, we have a variant of LIME in SoundLIME[10], which works on spectrograms or other audio-based features, treats them as image data only and then replicates the LIME framework. We use SoundLIME out of the black box models for our analysis. We employ both white-box and black-box XAI models to ensure comprehensive coverage of interpretability techniques, enabling a thorough comparison and analysis of the performance of our ARI model.
@@ III Dataset
The dataset comprises a total of 501 HCM songs, with each song being a polyphonic audio containing vocals of the main artist, percussion instruments like tabla, Pakhawaj, along with Manjeera in some audios, along with non-percussion instruments like Harmonium, Sarangi, Violin. These music recordings are sung in variety of Ragas and Talas with all of the songs being of varying lengths. The dataset includes 144 unique Ragas performed by 169 vocal artists. Harmonium is the most common instrument, present in a total of 372 songs, while some performances feature multiple string instruments. All of the songs include drone (tanpura) playing in the background of the performance, which provides the tonic for the whole music performance. The dataset includes performances in 20 different talas accompnying the vocals, with teentaal (121 instances) and ektaal (146 instances) being the most frequently used talas as shown in Fig. 1. Each song is available in the form of 30-second chunks, totaling 23,005 audio files, approximately 191.70 hours of music.
@@ III-B Recording Source and Quality
All recordings are in the form of mono tracks, sourced from the archives of Prasar Bharti (Akashwani, Delhi), which are not available online elsewhere. They feature performances by HCM ustaads (experts) and are mainly studio recordings or recordings from live shows, ensuring high audio quality. Most of the featured vocal, percussion, or string instrument artists are well-known in the HCM community. Some of the recordings feature speech at the start which includes introduction of the artists along with details about their performance including the Raga they will be performing. The speech segments range from a few seconds to about a minute or two in some recordings.
@@ III-C Metadata
The metadata contains details such as the names of vocal and instrumental artists, the instruments played, the event where the performance took place, the name of the song, and the number of 30-second audio files for each song. We also add to the metadata the Raga name, the tonic and the start and end times of music and speech for each of those songs, labeled manually by our annotators.
@@ III-D Manual Annotation Process
The labeling process for the Raga and Tonic labels is carried out by two expert vocalists, each with over 10 years of experience in HCM. The experts are provided with 501 complete audio files, each containing full performances, to label Raga names and tonic values. These labels are then applied to all the 30-second chunks derived from each file.
In many of these audio files, there is a narrator who announces the raga to be performed, ensuring the correctness of the raga label. Additionally, certain files come with pre-assigned Raga labels from Prasar Bharti. For these files, our experts review and verify the pre-assigned labels, half files each. For the remaining files without pre-assigned labels, the two experts independently label half of the files each. To further ensure consistency in labeling and minimize subjectivity, a third expert reviews all the Raga and Tonic labels throughout the dataset.
@@ IV Methodology
In this study, we use our curated dataset to extract chromagram features from the audios and train a CNN-LSTM model to perform multi-class classification for $N=12$ Raga classes, which represent the top classes with the most audio files in our dataset, ensuring sufficient examples for splitting the data into training, validation, and test sets. By limiting $N$ to just 12 classes, we aim to create a high performing classifier model, which is crucial for producing reliable results in our XAI analysis. Next, we employ Grad-CAM++ (GC) and SoundLIME (SL), two popular XAI models, for saliency visualization and compare them with annotations from music experts. We also train the same model on the whole dataset, but we do not use it for XAI analysis because of performance shortcomings. In this section, we explain the entire procedure.